# Towards a Dynamic Future with Adaptable Computing and Network Convergence (ACNC)

### v1

@@ III-B3 The E2E Orchestrator
The primary responsibility of the E2E orchestrator is to aggregate reduced states from both domains on a per-service basis, form a comprehensive system state denoted as , as depcited in Fig. 3, and store it in the memory bank. Upon securely storing this state in the memory bank, the orchestrator promptly communicates any significant discrepancies in the system state to PoAs via the network orchestrator, ensuring their active involvement in decision-making. Graph Edit Distance (GED) is utilized for measuring the minimum cost of transforming into through a sequence of edit operations. If the distance exceeds , triggering a change is considered. Simultaneously, is labeled with a context within a set of contexts of size . If the identified context is not active, PoAs are instructed to perform a context switch. ML-based techniques, such as Graph Convolutional Networks (GCNs), can be utilized for context detection. The context set can be predefined using historical data or expert knowledge and adjusted dynamically during system runtime. Further tasks involve optimizing rewards for PoAs to align with system-level objectives. Generative Adversarial Networks (GANs) can be employed to simultaneously train a generator and a discriminator, deriving rewards to attain optimal system states based on the target objectives. For example, it may involve reducing rewards for utilizing spare network links to conserve energy. Through hierarchical dimension reductions, context-based resource allocation, and differential state updating, minor oscillations (devoid of discernible impact) are disregarded, and adjustments in resource allocation are exclusively triggered for substantial variations with targeted consequences, enabling efficient control within the dynamically fluctuating environment of ACNC.
The E2E orchestrator’s final task involves reorganizing implemented instances and their respective E2E network paths from PoAs. ML-based techniques, such as Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) models, analyze historical/temporal data from the last system states to predict system feasibility in upcoming time slots. Upon detecting infeasibility, the E2E orchestrator employs a GNN-based RL agent to efficiently select computing nodes for migration or generation of instances and determine suitable E2E paths. Subsequently, the E2E orchestrator collaborates with the network orchestrator to inform PoAs of context changes and action sets, and with the computing orchestrator to facilitate new instance deployments. The CNC concept optimizes resource utilization across domains by considering the state of both domains, thereby assisting the orchestrator in this manner.
@@ III-B4 The Service Orchestrator
This layer assumes responsibility for overseeing the service lifecycle within the ACNC framework. Its initial function entails the reception and processing of requests originating from service providers, aiming to incorporate new services. These requests should include crucial service parameters, such as essential service instances, instance chaining maps, data models, and QoS/E requirements. Comprehensive records of these details are archived within the memory bank. The process of deploying new service instances across the infrastructure is orchestrated by the resource arrangement component of the E2E orcherstrator, which bases its decisions on the anticipated demand reported by the service provider. The service orchestrator’s additional responsibility is to systematically monitor all implemented services in an E2E fashion, maintaining a real-time, abstracted representation of their quality for subsequent reporting to the respective service provider. Beyond the specifics of registered services, the memory bank houses a repository of pre-trained agents, which play a vital role in efficient decision-making, especially during the system’s initial phases when there is no live data available for agent training. Furthermore, the memory bank includes simulation tools and digital twins to assess the system’s efficiency and determine potential adjustments. It also comprises training libraries like Torch or Tensor to support agent execution across various layers, and expert knowledge to fine-tune trained agents for optimized performance.
@@ III-C Integrating ACNC with SRv6
To facilitate the transmission of packets for each request session, ensuring they follow their selected network path to reach the selected service instances, the involvement of network devices along the path is essential. One viable solution for conveying this information is the utilization of the Segment Routing over IPv6 (SRv6) protocol. SRv6 operates on the principles of source routing, where the routing path is encoded within the packet header as a sequenced set of segments or path hops. Each network device along the list of segments, upon packet reception, removes its associated label from the list and directs the packet toward the network device indicated by the next label. This transmission occurs directly in cases where the list is defined in a hop-by-hop manner or via the shortest path if the path is divided into segments, with one network device designated per segment. SRv6 offers the advantage of precise traffic control without the necessity for supplementary protocols or path signaling, which results in a highly scalable solution.
@@ V-A Model Pre-Training: Digital Twin Integration
A promising avenue for future research in the ACNC framework involves exploring advanced methodologies for pre-training models and agents. Building models from scratch for each context, given the frequent and dynamic changes in the data plane and the diversity of involved technologies, may lead to suboptimal decision-making. The integration of digital twins, virtual counterparts mirroring the system and its behaviors, emerges as a promising strategy. By incorporating digital twins into the pre-training process and leveraging historical data for each service and context, models can effectively perform in real-world contexts from the initial time slot. Key steps include identifying system components, modeling their behavior, collecting relevant data, constructing request and traffic trends, and replicating the system’s behavior when all elements are integrated. A thorough investigation into these steps over the system components is essential to generating efficient pre-trained models for the ACNC framework.
@@ V-B Complexities in E2E ACNC Evaluation
In contrast to conventional communication systems, where E2E performance metrics are primarily influenced by predictable networking devices and protocols, evaluating system-wide performance in the ACNC framework poses a complex challenge. The dynamic allocation of numerous network devices and computing nodes within ACNC, coupled with the unpredictable nature of ML-based decision-making, complicates pinpointing precise sources of performance degradation. Therefore, diverse services or users may perceive performance variations differently. This underscores the need for an innovative approach to performance evaluation. Pioneering research initiatives aiming to establish a theoretical approach for measuring E2E performance in scenarios with multiple ML agents introduced to the data path will be crucial for the comprehensive utilization of ACNC.
@@ V-C ML Efficiency: Balancing Precision and Cost
In the process of constructing system states and in all optimization tasks of the E2E orchestrator, adjustable parameters play a crucial role in balancing precision and cost. Increasing the number of short-term state features ( and ), the number of reduced state indicators ( and ), the time window of long-term states ( and ), the number of possible contexts (), and the depth of the system state trajectory for feasibility check (), or decreasing the threshold of system state difference (), yields more detailed state data. While this enhances the precision of ML agents’ decisions, it necessitates increased storage for memory banks, higher network bandwidth and longer time for data exchange, and additional computing capacity and time for training and decision-making. Adjusting these parameters is a prospective research avenue to maximize precision in resource allocation while minimizing network and computing costs. Dynamically navigating the trade-offs between precision and cost in ML-based decision-making represents a pivotal research initiative with profound implications for future applications.

### v2

@@ III-B3 The E2E Orchestrator
The primary responsibility of the E2E orchestrator is to organize service instances and potential network paths for the next time slot. First, it aggregates reduced states from both domains to form a comprehensive system state, denoted as , which is stored in the memory bank (depicted in Figure 3). Since the network and computing orchestrators prepare the bindings of service instances with their respective resources, the E2E orchestrator concatenates these to generate a unified service-network-computing binding for all services, providing a system-wide view. Simultaneously, is labeled with a context from a set of contexts of size . To achieve this, we first process the state using a GNN, then apply Vector Quantization (VQ) to the GNN’s output to create embeddings for each state, grouping similar states into the same context based on the distance between the input state and the available contexts [12]. Each context corresponds to a cluster of similar states and is stored in the memory bank. Subsequently, the E2E orchestrator optimizes the reward functions of PoAs to align with system-level objectives—for example, selecting cheaper allocations when the objective is minimizing cost. When the accuracy of allocations in PoAs falls below a predetermined threshold, the Adaptive Resonance Theory (ART) is utilized to update the available set of contexts, which comprises two bidirectionally connected layers: an input layer (F1) and a recognition layer (F2) [13]. Upon presenting a state to F1, neurons in F2 are activated based on the similarity between the input and stored context weight vectors, potentially leading to the generation of new contexts as necessary. The implementation of ART within the framework of VQ facilitates the dynamic adjustment of the context set.
After preparing the system state and determining the current context, Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU) models are employed to predict the next system state by analyzing historical temporal data from the last system states [14]. These models incorporate recurrent connections and maintain an internal memory, enabling them to capture both short-term and long-term dependencies in the data for effective future state forecasting. A check is then performed on the predicted next state to detect any potential infeasibilities based on the requirements of the implemented services. Upon detecting infeasibility, the E2E orchestrator employs a RL agent to efficiently select new computing nodes for service instances. The agent uses GNN layers to process the system state represented as a graph-structured data. It is pre-trained to select a computing node for each service instance in each specific system state, optimizing for the system objective. Based on the selected computing nodes, potential E2E paths from each PoA to the chosen computing nodes are determined. The E2E orchestrator then collaborates with the network orchestrator to inform PoAs of the updated system state, context changes, as well as their action set, and with the computing orchestrator to facilitate new instance deployments.
@@ III-B4 The Service Orchestrator
This layer assumes responsibility for overseeing the service lifecycle within the ACNC framework. Its primary function involves receiving, processing, and storing requests from service providers, aiming to incorporate new services. These requests should include crucial service parameters, such as essential service instances, instance chaining maps, data models, and QoS/E requirements. The process of deploying new service instances across the infrastructure is orchestrated by the resource arrangement component of the E2E orcherstrator, which also includes the anticipated demand reported by the service provider. Additionally, it monitors all implemented services in an E2E fashion, maintaining a real-time, abstracted representation of their quality for subsequent reporting to the respective service provider. Beyond the specifics of registered services, the memory bank houses a repository of pre-trained agents, which play a vital role in efficient decision-making, especially during the system’s initial phases when there is no live data available for agent training. Furthermore, this includes simulation tools and digital twins to assess the system’s efficiency and determine potential adjustments.
@@ III-C Integrating ACNC with SRv6
To facilitate the transmission of packets for each request session, the involvement of network devices along the path is essential. One viable solution for conveying this information is the utilization of the Segment Routing over IPv6 (SRv6) protocol. SRv6 operates on the principles of source routing, where the routing path is encoded within the packet header as a sequenced set of segments or path hops. Each network device along the list of segments, upon packet reception, removes its associated label from the list and directs the packet toward the network device indicated by the next label. This transmission occurs directly in cases where the list is defined in a hop-by-hop manner or via the shortest path if the path is divided into segments, with one network device designated per segment. SRv6 offers the advantage of precise traffic control without the necessity for supplementary protocols or path signaling, which results in a highly scalable solution.
@@ V-A Model Pre-Training: Digital Twin Integration
A promising direction for future research in the ACNC framework is the exploration of advanced pre-training methodologies for models and agents. Building models from scratch for each context can lead to suboptimal decision-making due to the frequent changes in the data plane and the diversity of technologies involved. Integrating digital twins emerges as a valuable strategy. By incorporating digital twins into the pre-training process and leveraging historical data, models can perform effectively in real-world situations from the outset. Key steps include identifying system components, modeling their behavior, collecting relevant data, constructing request and traffic trends, and replicating the system’s behavior when all elements are integrated. A thorough investigation of these steps is essential for generating efficient pre-trained models within the ACNC framework.
@@ V-B Complexities in E2E ACNC Evaluation
In contrast to conventional communication systems, where E2E performance metrics are primarily influenced by predictable networking devices and protocols, evaluating system-wide performance in the ACNC framework poses a complex challenge. The dynamic allocation of numerous network devices and computing nodes within ACNC, coupled with the unpredictable nature of ML-based decision-making, complicates pinpointing precise sources of performance degradation. Therefore, diverse services or users may perceive performance variations differently. Pioneering research initiatives aiming to establish a theoretical approach for measuring E2E performance in scenarios with multiple ML agents introduced to the data path will be crucial for the comprehensive utilization of ACNC.
@@ V-C ML Efficiency: Balancing Precision and Cost
In the process of constructing system states and in all optimization tasks of the E2E orchestrator, adjustable parameters play a crucial role in balancing precision and cost. For instance, increasing the number of state features and indicators, as well as the history window size, provides more detailed state data. While this enhances the precision of ML agents’ decisions, it necessitates increased storage for memory banks, higher network bandwidth and longer time for vertical data exchange, and additional computing capacity and time for training and decision-making. Adjusting these parameters is a prospective research avenue to maximize precision in resource allocation while minimizing network and computing costs. Dynamically navigating the trade-offs between precision and cost in ML-based decision-making represents a pivotal research initiative with profound implications for future applications.